OpenStudy (anonymous):

The line through (2, -1) and parallel to x - 2y = 6.

OpenStudy (anonymous):

paralle lines have the same slope, so just solve for b like before

OpenStudy (anonymous):

(x-2)-2(y+1)=0 x-2y=4

OpenStudy (anonymous):

solve for y in x-2y=6 to find the right slope

OpenStudy (anonymous):

or just use standart form, which is way faster, :)